OpenStudy (anonymous):

implicit differentiation 9x^2+y^2=9

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

assuming you have to find \(\displaystyle\frac{\partial y}{\partial x}\) \[ 9x^2+y^2=9\\ \frac{\partial}{\partial x}(9x^2+y^2)=\frac{\partial}{\partial x}(9)\\ \implies\frac{\partial}{\partial x}(9x^2)+\frac{\partial}{\partial x}(y^2)=0\\ \implies9\frac{\partial}{\partial x}(x^2)+2y\frac{\partial y}{\partial x}=0\\ \implies9(2x)+2y\frac{\partial y}{\partial x}=0\\ \implies 2y\frac{\partial y}{\partial x}=-18x \]simplify further